One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
Answer by ptaylor(2052)

Let x=smaller number
Then x+12=the larger number
2(x+12)=twice the larger number
Now we are told that:
x+2(x+12)=39 get rid of parens
x+2x+24=39 subtract 24 from both sides
x+2x+24-24=39-24 collect like terms
3x=15 divide both sides by 3
x=5 -------------------------------smaller number
x+12=5+12=17----------------------------larger number
CK
5+2(17)=39
5+34=39
39=39
5+12=17
17=17

Hope this helps-------ptaylor